OpenStudy (anonymous):

i need help slope intercept form m =2 and (-1,5)

OpenStudy (precal):

y=mx+b m=2 x=-1 y=5 5=2(-1)+b 5=-2+b add 2 to both sides 7=b y=2x+7

OpenStudy (anonymous):

i dont get add 2 what 2 sides like 5+2= 7 just add that y +2

OpenStudy (precal):

5=-2+b +2 +2 7=b do you know how to solve equations?